### v1

# The Covariant Stone-von Neumann Theorem for Locally Compact Quantum Groups

Abstract.
The Stone-von Neumann Theorem is a fundamental result which unified the competing quantum-mechanical models of matrix mechanics and wave mechanics. It’s mechanism of proof ultimately involved the study of unitary group representations on a Hilbert space. In this article, we continue the broad generalization set out in [HI, modular], analyzing representations of locally compact quantum dynamical systems defined on Hilbert modules, of which the classical result is a special case. We introduce a pair of modular representations which subsume numerous models which appear in the literature, and for certain coactions $(\mathbb{G},A,\alpha)$ recover the multiplicity results of [modular]. As a corollary, we develop a new criterion for identifying strongly regular locally compact quantum groups, related to the study of their dynamics on elementary $C^{*}$-algebras.
1. Introduction
The fields of science have long benefited from interdisciplinary interactions, and the application of mathematics to these fields is no exception. The past century has arguably witnessed among the most influential discoveries in human history the discovery of quantum phenomena. With its foundation in mathematics and home in physics, the exploration of quantum mechanics has beyond any doubt cast us into the atomic era. Chief in the development of this knowledge were the independently developed models of quantum mechanics, namely Werner Heisenberg’s matrix mechanics and Erwin Schrödinger’s wave mechanics. Where matrix mechanics proposed a governance of subatomic phenomena by certain interacting linear transformations, wave mechanics distinguished itself by the application of partial differential equations. Remarkably, each model yielded the same experimental predictions, suggesting a unity of theory despite cosmetic differences. Using sophisticated new tools developed by Marshall Stone (which today serve as important rudiments in the study of operator algebras), John von Neumann successfully proved the formal mathematical equivalence of these two models in what is now called the Stone-von Neumann Theorem. Loosely stated, the theorem proves that any self-adjoint representation of Heisenberg’s linear operators is merely a multiple of the irreducible representation coming from solutions to Schrödinger’s wave equation. The essential strategy involved associating to a coupled pair of linear transformations two families of linear operators indexed by the topological group $\mathbb{R}^{n}$.
Von Neumann proceeded to study rings of operators, initiating the study of operator algebras which reaches as far in the field of mathematics as quantum mechanics does in physics. Indeed, operator algebraists have developed deep interactions with numerous “classical” mathematical fields, mirroring the drive to unify physics under one “standard model”. Throughout this enrichment of the field, a steady interest in Stone and von Neumann’s previous investigations remained, sustaining a perpetual refinement and generalization of ideas. Major influences include George Mackey and Marc Rieffel, with successive innovations from numerous others. Recently, Ismert and the second author provided a significant innovation, classifying representations of $C^{*}$-dynamical systems of abelian groups represented on Hilbert modules. Quigg and the first two authors expanded this classification to include dynamical systems of general (nonabelian) locally compact groups using divergent techniques from the parallel but distinct lenses of duality for $C^{*}$-actions and coactions.
From the lens of nonabelian duality, the divergence of technique is a necessary cost for the expansion beyond actions of an abelian group. However, over the past few decades the appealing new theory of locally compact quantum groups has emerged to unify these parallel developments under a common theory. In this article, we call upon this theory to integrate the work of Quigg and the first two authors, recovering numerous independent results inspired by Stone and von Neumann’s century-old work. Figuring prominently in this article is the emphasis on corepresentations of locally compact quantum groups, which are a valuable tool for quantum groups which appears implicitly in the coaction setting of [modular]. These objects, other conventions for locally compact quantum groups, and their crossed products are recorded in §2. Heisenberg representations of $C^{*}$-quantum group dynamical systems are defined in §3, where we also explore close relationships with prior evolutionary expressions. Together with these relationships, Theorem 3.9 serves to confirm the proposed definition. In §4, we introduce the Schrödinger representation of a quantum dynamical system, and prove our ultimate result, that every Heisenberg representation of a maximal coaction on an elementary $C^{*}$-algebra is a multiple of the Schrödinger representation. We conclude our investigation with an examination of conditions under which one expects a Stone-von Neumann-type result to hold, which includes results offering new sufficient conditions for maximality of a coaction as well as strong regularity for a locally compact quantum group.
Acknowledgments
The work of JK was partially supported by FWO grant 1246624N. as well as EPSRC grants EP/T03064X/1 and EP/T030992/1.
LHall recognizes the Zuckerman Institute, who supports him as a Zuckerman Postdoctoral Scholar in the 2023-2024 cohort.
MT was partially supported by NCN grant 2020/36/C/ST1/00082.
2. Preliminaries
In this section we record the conventions we will use throughout the article. All homomorphisms are assumed to be ***-preserving. An elementary $C^{*}$-algebra is a $C^{*}$-algebra $A$ which is isomorphic to the $C^{*}$-algebra of compact operators over some Hilbert space $H$.
2.1. Hilbert Modules
We refer to [lance, rw, buss] for details on Hilbert modules. Let $A$ be a $C^{*}$-algebra. A (right) Hilbert A-module is a complex vector space $X$, denoted $X_{A}$ when confusion seems possible, which is also a right $A$-module equipped with an $A$-valued inner product $\langle\cdot,\cdot\rangle_{A}\colon X\times X\to A$ satisfying various compatibility conditions. The inner product induces a natural norm on $X$, $\|x\|=\|\langle x,x\rangle_{A}\|_{A}^{1/2}$, and $X$ is required to be complete with respect to this norm; $X$ is called (right) full provided that the ideal $\text{span}\{\langle y,x\rangle\,|\,y,x\in X\}$ is dense in $A$. Standard examples include any $C^{*}$-algebra $A$, which is a Hilbert $A$-module with inner product $\langle b,a\rangle_{A}=b^{*}a$; and Hilbert $\mathbb{C}$-modules, which are Hilbert spaces with the convention that the ordinary inner product be conjugate-linear in the first variable.
The set of adjointable operators on $X$ is defined as
$\mathcal{L}(X)=\{T\colon X\to X|\text{ there is }T^{*}\colon X\to X\text{ suchthat }\langle T^{*}\eta,\xi\rangle_{A}=\langle\eta,T\xi\rangle_{A}\text{ for all }\xi,\eta\in X\}.$
$\mathcal{L}(X)$ is a $C^{*}$-algebra which generalizes bounded Hilbert-space operators. When $X=A$ is a $C^{*}$-algebra, we denote $\mathcal{L}(A)=M(A)$ and call $M(A)$ the multiplier algebra of $A$ — it is the universal unital $C^{*}$-algebra which contains $A$ as an essential ideal. A $B-A$ correspondence is a Hilbert $A$-module $X$ together with a nondegenerate homomorphism $B\to\mathcal{L}(X),$ by which we mean that $\overline{\operatorname{span}}\{b\cdot x|~{}b\in B\text{ and }x\in X\}=X$ (or equivalently $\{b\cdot x|~{}b\in B\text{ and }x\in X\}=X$ by the Cohen-Hewitt Factorization Theorem [hewitt]). A $B-A$ imprimitivity bimodule is a $B-A$ correspondence $X$ equipped with a $B$-valued inner product $\langle\cdot,\cdot\rangle_{B}\colon X\times X\to B$ which is left full and compatible with the $A$-valued inner product in the sense that $\langle\zeta,\eta\rangle_{B}\cdot\xi=\zeta\cdot\langle\eta,\xi\rangle_{A}$. For us, there are two imprimitivity bimodules of fundamental importance. The first is a Hilbert space $H$, which is a $\mathbb{K}(H)-\mathbb{C}$ imprimitivity bimodule. The left action is given by the identity map, and the left-inner-product structure is determined by the rule $\langle z,y\rangle_{\mathbb{K}(H)}=\theta_{z,y}$, where $\theta_{z,y}$ is the standard rank-one operator $\theta_{z,y}(x)=z\langle y,x\rangle_{\mathbb{C}}$ — compatibility of the inner products is by definition. The other natural imprimitivity bimodule is given by a $C^{*}$-algebra $A$, which is an $A-A$ imprimitivity bimodule. The left action is simply left multiplication, and the left-inner-product structure is given by $\langle c,b\rangle=cb^{*},$ thus $\langle c,b\rangle a=cb^{*}a=c\langle b,a\rangle.$ Fullness follows from an approximate-identity argument.
In [modular], Quigg and the first two authors completely classified the correspondences over elementary $C^{*}$-algebras. In the theorem below, one correspondence $Y$ is a multiple of another $X$ provided $Y$ is unitarily equivalent to a direct sum $\oplus_{S}X$.
Theorem 2.1.
[modular]*Theorem 3.5 Let $A$ be an elementary $C^{*}$-algebra and $X$ an $B-A$ imprimitivity bimodule. Then every $B-A$ correspondence $Y$ is multiple of $X$.
We assume familiarity with the spatial tensor product of $C^{*}$-algebras which will simply be denoted by $\otimes$. Given Hilbert modules $X_{A}$ and $Y_{B}$, one can define an external tensor product $X\otimes Y$, which is a Hilbert $A\otimes B$-module and contains the algebraic tensor product of $X$ and $Y$ as a dense subspace. Given $T\in\mathcal{L}(X)$ and $S\in\mathcal{L}(Y)$, then there is $T\otimes S\in\mathcal{L}(X\otimes Y)$ determined on elementary tensors in the obvious way. We will make extensive use of leg-numbering notation, which we briefly describe. For $T\in\mathcal{L}(X\otimes X),$ $T_{ij}\in\mathcal{L}(X\otimes X\otimes X)$ are defined for $1\leq i<j\leq 3$ by $T_{12}=T\otimes 1$, $T_{23}=1\otimes T$, and $T_{13}=(1\otimes\Sigma)T_{12}(1\otimes\Sigma)$, where $\Sigma\colon x\otimes y\mapsto y\otimes x$ is the flip operator on $X\otimes X$. For the flip map on the tensor product of two algebras $A\otimes B$, we will use symbol $\chi$.
2.2. The Millinery of Locally Compact Quantum Groups
Our main reference for locally compact quantum groups are [kv, buss]. Other standard references include [kustermansuniversal, venchilada, KVvn].
A locally compact quantum group $\mathbb{G}$ is a $C^{*}$-algebra $C_{0}(\mathbb{G})$ together with a nondegenerate ***-homomorphism $\Delta\colon C_{0}(\mathbb{G})\to M(C_{0}(\mathbb{G})\otimes C_{0}(\mathbb{G}))$ called comultiplication which satisfies the coassociativity and density conditions
$\begin{split}(\Delta\otimes\text{id})\Delta&=(\text{id}\otimes\Delta)\Delta,\text{ and}\\ C_{0}(\mathbb{G})=\overline{\operatorname{span}}\{(\omega\otimes\text{id})\Delta(x)|~{}\omega\in C_{0}(\mathbb{G})^{*},x\in C_{0}(\mathbb{G})\}&=\overline{\operatorname{span}}\{(\text{id}\otimes\omega)\Delta(x)|~{}\omega\in C_{0}(\mathbb{G})^{*},x\in C_{0}(\mathbb{G})\}.\end{split}$
Furthermore, by definition, there are two faithful, approximately KMS weights $\varphi,\psi$ on $C_{0}(\mathbb{G})$ which are respectively left- and right-invariant (for details see [kv, Definition 4.1]). One can equivalently work in the von Neumann-algebraic setting ([KVvn, Definition 1.1]), but we will stay in the language of noncommutative topology, i.e., $C^{*}$-algebras.
Associated with the left Haar weight $\varphi$ is the GNS Hilbert space, denoted $L^{2}(\mathbb{G})$, and faithful representation $\pi_{\varphi}\colon C_{0}(\mathbb{G})\to\mathbb{B}(L^{2}(\mathbb{G}))$. Since $\pi_{\varphi}$ is faithful, it is customary to consider $C_{0}(\mathbb{G})$ as acting on $L^{2}(\mathbb{G})$ and neglect writing $\pi_{\varphi}$, and we will follow this custom. One defines a unitary $\mathrm{W}^{\mathbb{G}}\in\mathbb{B}(L^{2}(\mathbb{G})\otimes L^{2}(\mathbb{G}))$ called the Kac-Takesaki operator. It is a multiplicative unitary, i.e., it satisfies the Pentagonal Relation
$\mathrm{W}^{\mathbb{G}}_{12}\mathrm{W}^{\mathbb{G}}_{13}\mathrm{W}^{\mathbb{G}}_{23}=\mathrm{W}^{\mathbb{G}}_{23}\mathrm{W}^{\mathbb{G}}_{12}.$
From $\mathrm{W}^{\mathbb{G}}$, one recovers the $C^{*}$-algebra $C_{0}(\mathbb{G})=\overline{\operatorname{span}}\{(\text{id}\otimes\omega)(\mathrm{W}^{\mathbb{G}})|~{}\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}\}$ and comultiplication $\Delta(a)=\mathrm{W}^{\mathbb{G}*}(1\otimes a)\mathrm{W}^{\mathbb{G}}\;(a\in C_{0}(\mathbb{G}))$. One can define the dual locally compact quantum group $\widehat{\mathbb{G}}$ — its $C^{*}$-algebra may be recognized as $C_{0}(\widehat{\mathbb{G}})=\overline{\operatorname{span}}\{(\omega\otimes\text{id})(\mathrm{W}^{\mathbb{G}})|~{}\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}\}$ and the comultiplication is given by $\widehat{\Delta}(a)=\mathrm{W}^{\widehat{\mathbb{G}}*}(1\otimes a)\mathrm{W}^{\widehat{\mathbb{G}}}\;(a\in C_{0}(\widehat{\mathbb{G}}))$, where $\mathrm{W}^{\widehat{\mathbb{G}}}=\chi(\mathrm{W}^{\mathbb{G}*})$. In fact, $\mathrm{W}^{\widehat{\mathbb{G}}}$ is the Kac-Takesaki operator of $\widehat{\mathbb{G}}$. The remaining features of $\widehat{\mathbb{G}}$ will be likewise decorated with hats. One can prove that in fact $\mathrm{W}^{\mathbb{G}}\in M(C_{0}(\mathbb{G})\otimes C_{0}(\widehat{\mathbb{G}}))$ and
$(\Delta\otimes\text{id})\mathrm{W}^{\mathbb{G}}=\mathrm{W}^{\mathbb{G}}_{13}\mathrm{W}^{\mathbb{G}}_{23},\qquad(\text{id}\otimes\widehat{\Delta})\mathrm{W}^{\mathbb{G}}=\mathrm{W}^{\mathbb{G}}_{13}\mathrm{W}^{\mathbb{G}}_{12}.$
The culmination of the theory is the recovery of generalized Pontryagin duality $\widehat{\widehat{\mathbb{G}}}=\mathbb{G}$.
Every locally compact group $G$ gives rise to two quantum groups. The first one is $\mathbb{G}$, with associated $C^{*}$-algebra $C_{0}(\mathbb{G})=C_{0}(G)$, comultiplication acting by $\Delta(f)(x,y)=f(xy)$ and weights $\varphi,\psi$ given by integration with respect to left (resp. right) Haar measure. One typically identifies $\mathbb{G}$ with $G$. The second one, the dual quantum group, is described by objects studied in abstract harmonic analysis: $C_{0}(\widehat{G})=C_{r}^{*}(G)$ is the reduced group $C^{*}$-algebra and comultiplication acts on generators via $\widehat{\Delta}(\lambda(g))=\lambda(g)\otimes\lambda(g)$. Both Haar weights of $\widehat{\mathbb{G}}$ are equal to the Plancherel weight. Nonclassical analogs include compact quantum groups (which entails $C_{0}(\mathbb{G})$ possess a unit) and discrete quantum groups (where $C_{0}(\mathbb{G})$ is a direct sum of matrix algebras).
Out of a quantum group $\mathbb{G}$, one builds yet another $C^{*}$-algebra, $C_{0}^{u}(\mathbb{G})$, which should be thought of as a universal version of $C_{0}(\mathbb{G})$. It comes with its own comultiplication: a non-degenerate ***-homomorphism $\Delta^{u}\colon C_{0}^{u}(\mathbb{G})\rightarrow M(C_{0}^{u}(\mathbb{G})\otimes C_{0}^{u}(\mathbb{G}))$ and the reducing map, which is a ***-epimorphism $\lambda\colon C^{u}_{0}(\mathbb{G})\rightarrow C_{0}(\mathbb{G})$ respecting comultiplications (i.e. $\Delta\circ\lambda=(\lambda\otimes\lambda)\circ\Delta^{u}$). Furthermore, it has a very useful universal property concerning corepresentations of $\widehat{\mathbb{G}}$ (see Section 2.3 and [kustermansuniversal]).
Let $J$ be the modular conjugation associated to the left Haar weight. The commutant of $\mathbb{G}$ is a locally compact quantum group $\mathbb{G}^{\textnormal{c}}$ with $C^{*}$-algebra $C_{0}(\mathbb{G}^{\textnormal{c}})=JC_{0}(\mathbb{G})J$ and comultiplication $\Delta^{\textnormal{c}}(a)=(J\otimes J)\Delta(JaJ)(J\otimes J)$ for $a\in C_{0}(\mathbb{G}^{\textnormal{c}})$. Another locally compact quantum group associated with $\mathbb{G}$ is the opposite of $\mathbb{G}$ with $C_{0}(\mathbb{G}^{\textnormal{op}})=C_{0}(\mathbb{G})$ and $\Delta^{\textnormal{op}}=\chi\circ\Delta$. This millinery of examples are related by various relations, namely
$\widehat{\mathbb{G}^{\textnormal{op}}}=\widehat{\mathbb{G}}^{\textnormal{c}}\qquad\widehat{\mathbb{G}^{\textnormal{c}}}=\widehat{\mathbb{G}}^{\textnormal{op}}\qquad\mathbb{G}^{\textnormal{op},\textnormal{c}}=\mathbb{G}^{\textnormal{c},\textnormal{op}}\cong_{\widehat{J}J}\mathbb{G};$
The final relation states that the opposite of the commutant of $\mathbb{G}$ is unitarily equivalent to $\mathbb{G}$ by the unitary $\widehat{J}J$ (see [KVvn, Section 4]). Each of the quantum groups we consider above admits universal companions, and we will decorate the respective reducing maps using the same notation, e.g. $\hat{\lambda}\colon C_{0}^{u}(\hat{\mathbb{G}})\to C_{0}(\hat{\mathbb{G}})$, and similarly with $\lambda^{\textnormal{op}}$ and $\lambda^{c}$.
The last definition introduced here is the notion of regularity. By definition, a locally compact quantum group $\mathbb{G}$ is regular if $\overline{\operatorname{span}}\{(\text{id}\otimes\omega)(\Sigma\mathrm{W}^{\mathbb{G}})|~{}\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}\}=\mathbb{K}(L^{2}(\mathbb{G}))$. One can show that $\mathbb{G}$ is regular if and only if $\widehat{\mathbb{G}}$ is regular. Furthermore, all compact or discrete quantum groups are regular as well as quantum groups with trivial scaling group (see [BaajSkandalis, venchilada] and references therein).
2.3. Corepresentations
Given a locally compact quantum group $\mathbb{G}$ and a Hilbert $A$-module $X$, a left corepresentation of $\mathbb{G}$ on $X$ is a unitary $u\in\mathcal{L}(C_{0}(\mathbb{G})\otimes X)$ which satisfies $(\Delta\otimes\text{id})(u)=u_{13}u_{23}$. Likewise, a right corepresentation is a unitary $v\in\mathcal{L}(X\otimes C_{0}(\mathbb{G}))$ which satisfies $(\text{id}\otimes\Delta)(v)=v_{12}v_{13}$. The assignment $u\mapsto u^{\textnormal{op}}=\chi(u)^{*}$ determines a bijective correspondence between left corepresentations of $\mathbb{G}$ and right corepresentations of $\mathbb{G}^{\textnormal{op}}$. The Kac-Takesaki operator is a distinguished left corepresentation of $\mathbb{G}$ on $L^{2}(\mathbb{G})$. In fact, $\mathrm{W}^{\mathbb{G}}\in M(C_{0}(\mathbb{G})\otimes C_{0}(\widehat{\mathbb{G}}))$, and regarded this way we call $\mathrm{W}^{\mathbb{G}}$ the left regular corepresentation of $\mathbb{G}$ (on $C_{0}(\widehat{\mathbb{G}})$). We also have the right regular corepresentation of $\mathbb{G}$ (which is indeed a right corepresentation corresponding to a particular left corepresentation of $\mathbb{G}^{\textnormal{op}}$), which is a unitary $\mathrm{V}^{\mathbb{G}}\in M(C_{0}(\widehat{\mathbb{G}}^{\textnormal{c}})\otimes C_{0}(\mathbb{G}))$ satisfying $(\text{id}\otimes\Delta)(\mathrm{V}^{\mathbb{G}})=\mathrm{V}^{\mathbb{G}}_{12}\mathrm{V}^{\mathbb{G}}_{13}$. It also implements the comultiplication, namely $\Delta(x)=\mathrm{V}^{\mathbb{G}}(x\otimes 1)\mathrm{V}^{\mathbb{G}*}$ for $x\in C_{0}(\mathbb{G})$.
[…]
2.4. Coaction Crossed Products
Let $A$ be a $C^{*}$-algebra and $\mathbb{G}$ a locally compact quantum group. A left coaction of $\mathbb{G}$ on $A$ is a nondegenerate homomorphism $\alpha\colon A\to M(C_{0}(\mathbb{G})\otimes A)$ which satisfies the following two conditions
$(\text{id}\otimes\alpha)\circ\alpha=(\Delta\otimes\text{id})\circ\alpha\quad\text{ and }\quad\overline{\text{span}}\,(C_{0}(\mathbb{G})\otimes 1)\alpha(A)=C_{0}(\mathbb{G})\otimes A.$
The second condition is variously referred to as continuity, coaction nondegeneracy, or the Podleś condition in the literature. The data $(\mathbb{G},A,\alpha)$ is also called a left coaction, or otherwise a left dynamical system. One analogously defines a right coaction of $\mathbb{G}$ on $A$ by swapping the positions of $A$ and $\mathbb{G}$ in the codomain of $\alpha$; as above, left coactions of $\mathbb{G}$ on $A$ correspond to right coactions of $\mathbb{G}^{\textnormal{op}}$ on $A$. Notice that the comultiplication of a locally compact quantum group is an example of a left (and right) coaction of $\mathbb{G}$ on itself. Another example is given by the (left) trivial coaction of $\mathbb{G}$ on $A$, denoted $(\mathbb{G},A,\text{triv})$. Here, $\text{triv}(a)=1\otimes a$.
Remark.
In the classical setting of locally compact groups, right coactions appear as the standard choice. In the quantum setting, however, left corepresentations and consequently left coactions appear to be more natural. Of course, in view of the comments above, the two notions are really more or less equivalent — the burden of working with right coactions is purely notational. Nevertheless, in this article, we will have use for both left and right coactions.
Given a left coaction $(\mathbb{G},A,\alpha)$ and any Hilbert $B$-module $X=X_{B}$, a covariant representation of $(\mathbb{G},A,\alpha)$ on $X_{B}$ is a pair of nondegenerate homomorphisms $(\mu,\pi)$, with $\mu\colon C_{0}^{u}(\widehat{\mathbb{G}})\to\mathcal{L}(X_{B})$ and $\pi\colon A\to\mathcal{L}(X_{B})$, subject to the condition that
$(\text{id}\otimes\pi)\circ\alpha(a)=(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes\pi(a))(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$
for all $a\in A$. Notice that the right hand side is conjugation by the left unitary corepresentation $(\text{id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$. Elsewhere in the literature, a corepresentation is regarded as a pair comprising a homomorphism of the coefficient algebra and a left corepresentation on the common Hilbert module $X$. Given a right coaction $(A,\mathbb{G},\alpha)$, a covariant representation on $X=X_{B}$ is a pair $(\pi,m)$ with $\pi\colon A\to\mathcal{L}(X_{B})$ and $m\colon C_{0}^{u}(\widehat{\mathbb{G}}^{\textnormal{c}})\to\mathcal{L}(X_{B})$ subject to
$(\pi\otimes\text{id})\circ\alpha(a)=(m\otimes\text{id})(\mathds{V}^{\mathbb{G}})(\pi(a)\otimes 1)(m\otimes\text{id})(\mathds{V}^{\mathbb{G}})^{*}$
An important example of a covariant representation for $(\mathbb{G},A,\alpha)$ is given on the Hilbert module $L^{2}(\mathbb{G})\otimes A$ by the pair $(\hat{\lambda}\otimes 1,\alpha)$ – it is called the the regular representation of the coaction.
For the left coaction $(\mathbb{G},A,\alpha)$, the full crossed product is a $C^{*}$-algebra $\mathbb{G}\ltimes A$ together with a covariant representation
$j_{\mathbb{G}}\colon C_{0}^{u}(\widehat{\mathbb{G}})\to M(\mathbb{G}\ltimes A),\qquad j_{A}\colon A\to M(\mathbb{G}\ltimes A)$
which is universal in the following sense: whenever $(\mu,\pi)$ is a covariant representation of $(\mathbb{G},A,\alpha)$ on $X$, there is a unique nondegenerate homomorphism $\mu\ltimes\pi\colon\mathbb{G}\ltimes A\to\mathcal{L}(X)$, called the integrated form (of the covariant representation), so that
$\mu=(\mu\ltimes\pi)\circ j_{\mathbb{G}},\qquad\pi=(\mu\ltimes\pi)\circ j_{A}.$
On account of the above identifications, one may also disintegrate an integrated form $\mu\ltimes\pi$ to recover the original covariant representation by precomposing by the universal covariant representations. The full crossed product exists, is unique up to isomorphism, and moreover (see [venchilada, vergnioux])
(1) $\mathbb{G}\ltimes A=\overline{\rm span}\{j_{A}(a)j_{\mathbb{G}}(h)~{}|~{}a\in A,h\in C_{0}^{u}(\widehat{\mathbb{G}})\}.$
For example, the crossed product for the trivial coaction $(\mathbb{G},A,\text{triv})$ is the maximal tensor product $C_{0}^{u}(\widehat{\mathbb{G}})\otimes_{\textnormal{max}}A$. For other fundamental examples, one constructs the full crossed product $\mathbb{G}\ltimes C_{0}(\mathbb{G})$ for the coaction $(\mathbb{G},C_{0}(\mathbb{G}),\Delta_{\mathbb{G}})$, and the reduced crossed product $\mathbb{G}\ltimes_{r}C_{0}(\mathbb{G})=(\hat{\lambda}\otimes 1\ltimes\Delta_{\mathbb{G}})(\mathbb{G}\ltimes C_{0}(\mathbb{G}))$. A locally compact quantum group $\mathbb{G}$ is regular if and only if $\mathbb{G}\ltimes_{r}C_{0}(\mathbb{G})\cong\mathbb{K}(L^{2}(\mathbb{G}))$. Similarly, $\mathbb{G}$ is called strongly regular provided that $\mathbb{G}\ltimes C_{0}(\mathbb{G})\cong\mathbb{K}(L^{2}(\mathbb{G}))$ ([venchilada, Definition 2.11]).
Let $(\mathbb{G},A,\alpha)$ be again an arbitrary left coaction. There is a left coaction $\widehat{\alpha}^{\rm op}$ of $\widehat{\mathbb{G}}^{\rm op}$ on $\mathbb{G}\ltimes A$, called the dual coaction to $\alpha$, uniquely characterized by
(2) $\widehat{\alpha}^{\rm op}(j_{\mathbb{G}}(h))=(\hat{\lambda}\otimes j_{\mathbb{G}})\Delta^{u}_{\widehat{\mathbb{G}}^{\rm op}}(h),\quad h\in C^{u}_{0}(\widehat{\mathbb{G}}),\qquad\widehat{\alpha}^{\rm op}(j_{A}(a))=1\otimes j_{A}(a),\quad a\in A.$
Indeed, one may check that $((\hat{\lambda}\otimes j_{\mathbb{G}})\Delta_{\widehat{\mathbb{G}}^{\rm op}}^{u},1\otimes j_{A})$ forms a covariant representation of $(\mathbb{G},A,\alpha)$ on $C_{0}(\widehat{\mathbb{G}}^{\rm op})\otimes(\mathbb{G}\ltimes A)$. Then, by the universal property of $\mathbb{G}\ltimes A$, we obtain a non-degenerate ***-homomorphism $\widehat{\alpha}^{\rm op}=(\hat{\lambda}\otimes j_{\mathbb{G}})\Delta^{u}_{\widehat{\mathbb{G}}^{\rm op}}\ltimes(1\otimes j_{A})\colon\mathbb{G}\ltimes A\rightarrow M(C_{0}(\widehat{\mathbb{G}}^{\rm op})\otimes(\mathbb{G}\ltimes A))$ acting via (2).
Let us end this section with a definition: for a coaction $(\mathbb{G},A,\alpha)$ with $\mathbb{G}$ regular, there is a canonical surjection $\widehat{\mathbb{G}}^{\rm op}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))\otimes A,$ which is given by the integrated form of the covariant representation $(\lambda^{c}\otimes 1,(\hat{\lambda}\otimes 1)\ltimes\alpha)$; the coaction is called maximal provided this surjection is an isomorphism.
3. Heisenberg Representations
Modular representations for group actions on operator algebras already appear in [HI] and [modular], emphasizing pointwise notions of covariant representations. Representations of coactions lack such a pointwise notions, and require a distinct but parallel theory, which is more appropriate from the lens of locally compact quantum groups. Below, we offer a definition of Heisenberg representations which extends the acceptable dynamics and unifies the classical setting under the common theory of corepresentations.
Definition 3.1.
Let $(\mathbb{G},A,\alpha)$ be a left coaction. A Heisenberg representation $(X,\rho^{c},\widehat{\rho},\pi)$ of $(\mathbb{G},A,\alpha)$ consists of a Hilbert $A$-module $X$ and non-degenerate ***-homomorphisms
$\rho^{c}\colon C_{0}^{u}(\mathbb{G}^{c})\longrightarrow\mathcal{L}(X),\qquad\widehat{\rho}\colon C_{0}^{u}(\widehat{\mathbb{G}})\longrightarrow\mathcal{L}(X),\qquad\pi\colon A\longrightarrow\mathcal{L}(X),$
such that
  1. (R1)
    $(\widehat{\rho},\pi)$ is a covariant representation of $(\mathbb{G},A,\alpha)$ on $X$,
  2. (R2)
    $\mathrm{W}_{13}^{\mathbb{G}}({\rm id}\otimes\widehat{\rho}\,)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}=(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}({\rm id}\otimes\widehat{\rho}\,)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}$,
  3. (R3)
    The images of $\pi$ and $\rho^{c}$ commute.
One certainly wishes to translate between the definition offered above and prior literary iterants. The next two propositions accomplish this, followed by a brief investigation of the relationships with this article’s predecessors.
Proposition 3.2.
Let $\mathbb{G}$ be a locally compact quantum group, let $X$ be a right Hilbert module over a C*-algebra $A$, and let
$\rho^{c}\colon C_{0}^{u}(\mathbb{G}^{c})\longrightarrow\mathcal{L}(X),\qquad\widehat{\rho}\colon C_{0}^{u}(\widehat{\mathbb{G}})\longrightarrow\mathcal{L}(X),$
be non-degenerate ***-homomorphisms. The following are equivalent:
  1. (R2)
    $\mathrm{W}_{13}^{\mathbb{G}}({\rm id}\otimes\widehat{\rho}\,)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}=(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}({\rm id}\otimes\widehat{\rho}\,)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}$.
  2. (R2’)
    $(\widehat{\rho},\rho^{c})$ is a covariant representation of the right coaction $(C^{u}_{0}(\widehat{\mathbb{G}}),\widehat{\mathbb{G}},({\rm id}\otimes\hat{\lambda})\Delta_{\widehat{\mathbb{G}}}^{u})$ on $X$.
Proof.
To prove that (R2) implies (R2’), observe that
$({\rm id}\otimes\widehat{\rho}\otimes\hat{\lambda})({\rm id}\otimes\Delta^{u}_{\widehat{\mathbb{G}}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$ $=\mathrm{W}_{13}^{\mathbb{G}}({\rm id}\otimes\widehat{\rho}\,)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}$
$=(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}({\rm id}\otimes\widehat{\rho}\,)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}^{*}\,.$
Take $\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}$. After applying $(\omega\otimes{\rm id}\otimes{\rm id})$ to both sides of the above equation, we obtain
$(\widehat{\rho}\otimes\hat{\lambda})\Delta^{u}_{\widehat{\mathbb{G}}}((\omega\otimes{\rm id})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}}))=(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\widehat{\rho}((\omega\otimes{\rm id})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}}))\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})^{*}\,.$
Since the elements $(\omega\otimes{\rm id})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$ span a dense subspace of $C^{u}_{0}(\widehat{\mathbb{G}})$, the conclusion follows. The preceding equation defines a covariant representation of $(\widehat{\rho},\rho^{c})$, and the above computation can be easily reversed to yield the remaining implication. ∎
Proposition 3.3.
Let $\mathbb{G}$ be a locally compact quantum group, let $X$ be a right Hilbert module over a C*-algebra $A$, and let
$\pi\colon A\longrightarrow\mathcal{L}(X),\qquad\rho^{c}\colon C_{0}^{u}(\mathbb{G}^{c})\longrightarrow\mathcal{L}(X),$
be non-degenerate ***-homomorphisms. The following are equivalent:
  1. (R3)
    The images of $\pi$ and $\rho^{c}$ commute.
  2. (R3’)
    $(\pi,\rho^{c})$ is a covariant representation of the right coaction $(A,\widehat{\mathbb{G}},{\rm triv})$ on $X$.
Here, ${\rm triv}\colon A\rightarrow M(A\otimes C_{0}(\widehat{\mathbb{G}}))$ denotes the right trivial coaction given by ${\rm triv}(a)=a\otimes 1$.
Proof.
Assume that (R3) holds. To prove (R3’), we have to show that
$(\pi\otimes{\rm id}){\rm triv}(a)=(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\pi(a)\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})^{*}$
for all $a\in A$, which is equivalent to
$(\pi(a)\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})=(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\pi(a)\otimes 1).$
Let $\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}$. Then
$({\rm id}\otimes\omega)((\pi(a)\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}}))$ $=\pi(a)\rho^{c}(({\rm id}\otimes\omega)(\mathds{V}^{\widehat{\mathbb{G}}}))$
$=\rho^{c}(({\rm id}\otimes\omega)(\mathds{V}^{\widehat{\mathbb{G}}}))\pi(a)$
$=({\rm id}\otimes\omega)((\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\pi(a)\otimes 1)).$
Since the above equation holds for any $\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}$, the result follows. Again reversing the above computation yields the converse claim. ∎
We are now prepared to apply Definition 3.1 and recover the Heisenberg representations promoted in numerous historical contexts.
Example 3.4.
Consider a Heisenberg $(\mathbb{G},A,\alpha)$-representation $(X,\rho^{c},\hat{\rho},\pi)$ comprising $A=\mathbb{C}$, $X=H$ a Hilbert space, and $\mathbb{G}=G$ a classical abelian locally compact group. Then $\rho^{c}$ and $\widehat{\rho}$ correspond to unitary representations $\widehat{U}$ and $U$ of $\widehat{G}$ and $G$ on $H$ respectively. The conditions (R1) and (R3) are automatically satisfied. Let ${\rm ev}_{g}\colon C_{0}(G)\to\mathbb{C}$ and ${\rm ev}_{\chi}\colon C_{0}(\widehat{G})\to\mathbb{C}$ be the evaluation functionals at points $g\in G$ and $\chi\in\widehat{G}$ respectively. After applying $({\rm ev}_{g}\otimes{\rm id}\otimes{\rm ev}_{\chi})$ to the defining equation of condition (R2), we obtain
$\chi(g)U(g)\widehat{U}(\chi)=\widehat{U}(\chi)U(g),$
which is the Mackey’s generalization of the Weyl commutation relation [Mackey]*Theorem 1.
Example 3.5.
Rieffel refined and generalized Mackey’s pioneering work on induced representations of locally compact groups to the setting of of $C^{*}$-representation theory (see the introduction to [rinduced] for a nice history), which recovers the Mackey-Stone-von Neumann Theorem treated in Example 3.4 as a special case of Mackey’s Imprimitivity Theorem (follow the references in [Rieffel]*Paragraph Three). In the latter article, Rieffel defines Heisenberg $G$-modules for an arbitrary locally compact group $G$. These are exactly our Heisenberg $(\mathbb{G},A,\alpha)$-representations for $A=\mathbb{C}$ and $\mathbb{G}$ the dual to a classical locally compact group. The discerning will notice that the integrand in [Rieffel]*Equation (3) has striking similarities to the application of the left regular corepresentation interpreted in the classical setting, an antecedent to our current work.
Example 3.6.
If we take $\mathbb{G}$ to be a classical abelian locally compact group $G$ and $A$ to be the $C^{*}$-algebra of compact operators $\mathbb{K}$ in Definition 3.1, then we recover the definition of the Heisenberg representation given by Huang and Ismert [HI]. Indeed, they explicitly assume (R1), (R3’), and the Mackey–Weyl commutation relation, which is equivalent to (R2) (see the previous remark).
Example 3.7.
In [modular], the authors consider parallel generalizations of the Heisenberg representation using tools of nonabelian duality. In each case, the expressions differ in whether $\mathbb{G}$ is classical (hence derived from a bonafide locally compact group) or universal and dual to classical (i.e the full $C^{*}$-algebra of a locally compact group). Correcting for the cosmetics of left and right (co)actions, for classical $\mathbb{G}$ [modular]*Definition 4.5 defines a Heisenberg action-modular representation in accordance with our conditions (R1), (R2’), and (R3). Likewise, [modular]*Definition 5.1 defines a Heisenberg coaction-modular representation in exactly the same way. Our present work is a proper generalization in that there are numerous locally compact quantum groups which are neither classical nor dual to classical.
Example 3.8.
In [mrw], the authors introduced the notion of a Heisenberg pair which Roy and Timmermann later generalized in [RT18]. Meyer, Roy, and Woronowicz define Heisenberg pairs in the context of locally compact quantum groups derived from manageable multiplicative unitaries (see, e.g. [Wor96]) and are closely related to our Heisenberg representations in the sense of Kustermans and Vaes. Compare for instance the parallels between our Proposition 3.9 and [RT18, Theorem 6.2]. It is known that Kustermans and Vaes’s description of locally compact quantum groups subsume the Woronowicz framework.
We end this section with a theorem which relates our Heisenberg modular representations to dynamics of the crossed product. Much as full crossed products completely capture the dynamics of the associated system, our Heisenberg representations retain the data involved in crossed product constructions.
Proposition 3.9.
Let $(\mathbb{G},A,\alpha)$ be a left coaction and $X$ a Hilbert $A$-module. There is a one-to-one correspondence between Heisenberg representations of $(\mathbb{G},A,\alpha)$ on $X$ and covariant representations of the left coaction $(\widehat{\mathbb{G}}^{\rm op},\mathbb{G}\ltimes A,\widehat{\alpha}^{\rm op})$ on $X$.
Proof.
Let $(X,\rho^{c},\widehat{\rho},\pi)$ be a Heisenberg representation. Define
$\mu:=\rho^{c},\qquad\Pi:=\widehat{\rho}\ltimes\pi.$
Since $(\widehat{\rho},\pi)$ is a covariant representation of $(\mathbb{G},A,\alpha)$ on $X$ by (R1), $\Pi$ is well defined. We have to prove that
(3) $({\rm id}\otimes\Pi)\widehat{\alpha}^{\rm op}(x)=({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})^{*}(1\otimes\Pi(x))({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})$
for all $x\in\mathbb{G}\ltimes A$. Using (1), it suffices to prove the above formula for the elements of the form $j_{A}(a)$, $a\in A$, and $j_{\mathbb{G}}(h)$, $h\in C^{u}_{0}(\widehat{\mathbb{G}})$. Observe that
$({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})({\rm id}\otimes\Pi)\widehat{\alpha}^{\rm op}(j_{A}(a))=(\text{id}\otimes\mu)({\text{\reflectbox{$\mathds{W}$}}\>\!}^{\widehat{\mathbb{G}}^{\rm op}})(1\otimes\Pi(j_{A}(a)))$
$=({\rm id}\otimes\rho^{c})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})(1\otimes\pi(a))=(1\otimes\pi(a))({\rm id}\otimes\rho^{c})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})$
$=(1\otimes\Pi(j_{A}(a)))({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}}).$
Here, we used the condition (R3). Similarly, we have that
$({\rm id}\otimes\Pi)\widehat{\alpha}^{\rm op}(j_{\mathbb{G}}(h))=({\rm id}\otimes\Pi)(\hat{\lambda}\otimes j_{\mathbb{G}})\Delta_{\widehat{\mathbb{G}}^{\rm op}}^{u}(h)=({\rm id}\otimes\widehat{\rho})(\hat{\lambda}\otimes{\rm id})\Delta_{\widehat{\mathbb{G}}^{\rm op}}^{u}(h)$
$=\chi((\widehat{\rho}\otimes{\rm id})({\rm id}\otimes\hat{\lambda})\Delta_{\widehat{\mathbb{G}}}^{u}(h))=\chi((\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\widehat{\rho}(h)\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})^{*})$
$=({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})^{*}(1\otimes\Pi(j_{\mathbb{G}}(h)))({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}}).$
Here we used the condition (R2’) and the formula $\chi(\mathds{V}^{\widehat{\mathbb{G}}})^{*}=\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}}$. We have proved equality (3) for generators, which ends proof in one direction.
Now let $(\mu,\Pi)$ be a covariant representation of $(\widehat{\mathbb{G}}^{\rm op},\mathbb{G}\ltimes A,\widehat{\alpha}^{\rm op})$ on $X$. Define
$\rho^{c}:=\mu,\qquad\widehat{\rho}:=\Pi\circ j_{\mathbb{G}},\qquad\pi:=\Pi\circ j_{A}\,.$
(R1) We have to prove that
$({\rm id}\otimes\pi)\alpha(a)=({\rm id}\otimes\widehat{\rho})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes\pi(a))({\rm id}\otimes\widehat{\rho})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$
for all $a\in A$. Since the above equation is equivalent to
$({\rm id}\otimes\Pi)(({\rm id}\otimes j_{A})\alpha(a))=({\rm id}\otimes\Pi)(({\rm id}\otimes j_{\mathbb{G}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes j_{A}(a))({\rm id}\otimes j_{\mathbb{G}})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})),$
the claim follows from the fact that $(j_{\mathbb{G}},j_{A})$ is a covariant representation of $(\mathbb{G},A,\alpha)$ on $\mathbb{G}\ltimes A$.
(R2’) We reverse one of the above calculations and obtain
$(\widehat{\rho}\otimes{\rm id})({\rm id}\otimes\hat{\lambda})\Delta_{\widehat{\mathbb{G}}}^{u}(h)$ $=\chi(({\rm id}\otimes\Pi)\widehat{\alpha}^{\rm op}(j_{\mathbb{G}}(h)))$
$=\chi(({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})^{*}(1\otimes\Pi(j_{\mathbb{G}}(h)))({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}}))$
$=(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})(\widehat{\rho}(h)\otimes 1)(\rho^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})^{*}.$
(R3) We have that
$({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})({\rm id}\otimes\Pi)\widehat{\alpha}^{\rm op}(j_{A}(a))=(1\otimes\Pi(j_{A}(a)))({\rm id}\otimes\mu)(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}}).$
Take $\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}$. After applying $(\omega\otimes{\rm id})$ to the above equation and using the definitions of $\rho^{c}$ and $\pi$, we find
$\rho^{c}((\omega\otimes{\rm id})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}}))\pi(a)=\pi(a)\rho^{c}((\omega\otimes{\rm id})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})).$
Since the elements $(\omega\otimes{\rm id})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\widehat{\mathbb{G}}^{\rm op}})$ span a dense subspace of $C^{u}_{0}(\mathbb{G}^{c})$, we are done. ∎
4. The Covariant Stone-von Neumann Theorem
We now introduce the Schrödinger representation of a dynamical system $(\mathbb{G},A,\alpha)$, which may be profitably considered as a fundamental example of a representation for the dynamics.
Definition 4.1.
Let $(\mathbb{G},A,\alpha)$ be a left coaction of $\mathbb{G}$ on $A$. A Schrödinger representation is a quadruple $(L^{2}(\mathbb{G})\otimes A,\lambda^{c}\otimes 1,\hat{\lambda}\otimes 1,\alpha)$, where
$\lambda^{c}\colon C^{u}_{0}(\mathbb{G}^{c})\longrightarrow C_{0}(\mathbb{G}^{c}),\qquad\hat{\lambda}\colon C^{u}_{0}(\widehat{\mathbb{G}})\longrightarrow C_{0}(\widehat{\mathbb{G}}),$
are the canonical surjective ***-homomorphisms.
Proposition 4.2.
The Schrödinger representation is a Heisenberg representation.
Proof.
(R1) We have to prove that $(\hat{\lambda}\otimes 1,\alpha)$ is a covariant representation of $(\mathbb{G},A,\alpha)$ on $L^{2}(\mathbb{G})\otimes A$, i.e.
$({\rm id}\otimes\alpha)\alpha(a)=({\rm id}\otimes(\hat{\lambda}\otimes 1))(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes\alpha(a))({\rm id}\otimes(\hat{\lambda}\otimes 1))(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}}).$
Note that the LHS equals $(\Delta_{\mathbb{G}}\otimes\text{id})\alpha(a)$ due to the covariance condition. We compute the RHS:
$({\rm id}\otimes(\hat{\lambda}\otimes 1))(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})^{*}(1\otimes\alpha(a))({\rm id}\otimes(\hat{\lambda}\otimes 1))(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})$ $=(\mathrm{W}^{\mathbb{G}}_{12})^{*}(1\otimes\alpha(a))\mathrm{W}^{\mathbb{G}}_{12}=(\Delta_{\mathbb{G}}\otimes\text{id})\alpha(a).$
Here, we used the fact that $\mathrm{W}^{\mathbb{G}}$ implements $\Delta_{\mathbb{G}}$.
(R2) It suffices to prove that
$\mathrm{W}_{13}^{\mathbb{G}}({\rm id}\otimes\hat{\lambda})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}(\lambda^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}=(\lambda^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}({\rm id}\otimes\hat{\lambda})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}\,,$
which follows from direct calculation
$\mathrm{W}_{13}^{\mathbb{G}}({\rm id}\otimes\hat{\lambda})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}(\lambda^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}=\mathrm{W}^{\mathbb{G}}_{13}\mathrm{W}^{\mathbb{G}}_{12}\mathrm{V}^{\widehat{\mathbb{G}}}_{23}=({\rm id}\otimes\Delta_{\widehat{\mathbb{G}}})(\mathrm{W}^{\mathbb{G}})\mathrm{V}^{\widehat{\mathbb{G}}}_{23}$
$=\mathrm{V}^{\widehat{\mathbb{G}}}_{23}\mathrm{W}^{\mathbb{G}}_{12}(\mathrm{V}^{\widehat{\mathbb{G}}}_{23})^{*}\mathrm{V}^{\widehat{\mathbb{G}}}_{23}=\mathrm{V}^{\widehat{\mathbb{G}}}_{23}\mathrm{W}^{\mathbb{G}}_{12}=(\lambda^{c}\otimes{\rm id})(\mathds{V}^{\widehat{\mathbb{G}}})_{23}({\rm id}\otimes\hat{\lambda})(\text{\reflectbox{$\mathds{W}$}}\>\!^{\mathbb{G}})_{12}\,.$
(R3) The images of $\alpha$ and $\lambda^{c}\otimes 1$ commute, since $\alpha(A)\subseteq M(C_{0}(\mathbb{G})\otimes A)$ and $(\lambda^{c}\otimes 1)(C_{0}^{u}(\mathbb{G}^{c}))\subseteq C_{0}(\mathbb{G}^{c})\otimes 1$. ∎
Using the machinery we have gathered so far, we can prove the main result of our work.
Theorem 4.3.
Let $\left(\mathbb{G},A,\alpha\right)$ be a left coaction, where $A$ is an elementary $C^{*}$-algebra, $\mathbb{G}$ is a regular locally compact quantum group, and $\alpha$ is a maximal coaction. Then every Heisenberg $\left(\mathbb{G},A,\alpha\right)$-representation is a multiple of the Schrödinger $\left(\mathbb{G},A,\alpha\right)$-representation.
Proof.
Consider a Heisenberg $(\mathbb{G},A,\alpha)$-representation $(X,\rho^{\textnormal{c}},\widehat{\rho},\pi)$ of $(\mathbb{G},A,\alpha)$. Then Proposition 3.9 yields a covariant representation $(\rho^{\textnormal{c}},\widehat{\rho}\ltimes\pi)$ for $(\widehat{\mathbb{G}}^{\rm op},\mathbb{G}\ltimes A,\widehat{\alpha}^{\rm op})$. In turn, the covariant representation $(\rho^{\textnormal{c}},\widehat{\rho}\ltimes\pi)$ determines a nondegenerate representation of the bicrossed product
$\rho^{\textnormal{c}}\ltimes\widehat{\rho}\ltimes\pi\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\to\mathcal{L}(X).$
Since $\alpha$ is maximal, we have an isomorphism $(\lambda^{c}\otimes 1)\ltimes(\hat{\lambda}\otimes 1)\ltimes\alpha\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\to\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$. Abbreviating the isomorphism by $\Phi$, we witness $X$ is an $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A-A$ correspondence with left action coming from the ***-homomorphism $(\rho^{c}\ltimes\hat{\rho}\ltimes\pi)\circ\Phi^{-1}$.
Applying this analysis to the Schrödinger representation $(L^{2}(\mathbb{G})\otimes A,\lambda^{c}\otimes 1,\hat{\lambda}\otimes 1,\alpha),$ we can also see that the correspondence $L^{2}(\mathbb{G})\otimes A$ is an $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A-A$ correspondence. It is in fact an imprimitivity bimodule. To see this, notice that the map $\Phi$ in the previous paragraph agrees with the bicrossed product ***-homomorphism associated to the Schrödinger representation, so that the left action of $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$ is given by the identity. Since $A$ and $\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$ are both simple, the fullness of $L^{2}(\mathbb{G})\otimes A$ is clear, and for the compatibility of the inner products, we compute for elementary tensors
$\langle h\otimes c,g\otimes b\rangle\cdot f\otimes a$ $=(\langle h,g\rangle\otimes\langle c,b\rangle)\cdot f\otimes a=(\theta_{h,g}\otimes cb^{*})\cdot f\otimes a=h\langle g,f\rangle\otimes cb^{*}a$
$=h\otimes c\cdot(\langle g,f\rangle\otimes\langle b,a\rangle)=h\otimes c(\langle g\otimes b,f\otimes a\rangle).$
Thus we may apply Proposition 2.1 to deduce that $X$ is a multiple of $L^{2}(\mathbb{G})\otimes A$. We suppress the unitary which implements the equivalence $X\cong\bigoplus_{S}L^{2}(\mathbb{G})\otimes A$, and proceed by identifying these two modules.
To finish, disintegrate the integrated form
$\rho^{\textnormal{c}}\ltimes\widehat{\rho}\ltimes\pi\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\to\mathcal{L}(\bigoplus_{S}L^{2}(\mathbb{G})\otimes A),$
recovering the covariant representation $(\rho^{\textnormal{c}},\hat{\rho}\ltimes\pi)$ of $(\widehat{\mathbb{G}}^{\textnormal{op}},\mathbb{G}\ltimes A,\hat{\alpha}^{\textnormal{op}}).$ Applying Proposition 3.9, we recover the Heisenberg $(\mathbb{G},A,\alpha)$-representation $(\oplus_{S}\big{(}L^{2}(\mathbb{G})\otimes A\big{)},\rho^{\textnormal{c}},\hat{\rho},\pi).$ Since each summand is invariant under the ***-representations, this concludes the proof. ∎
In light of Theorem 4.3, we introduce a new property for a left coaction $(\mathbb{G},A,\alpha)$, which is a quantum-dynamical analog of the behavior described in [Rieffel]*Theorem and [wcrossed]*Theorem 4.29.
Definition 4.4.
A dynamical system $(\mathbb{G},A,\alpha)$ is called Stone-von Neumann unique provided that every Heisenberg representation is a multiple of the Schrödinger representation.
By Theorem 4.3 if $\mathbb{G}$ is regular, $A$ is elementary and $\alpha$ is maximal, then $(\mathbb{G},A,\alpha)$ is Stone-von Neumann unique. In the other direction, given a dynamical system $(\mathbb{G},A,\alpha)$ which possesses Stone-von Neumann uniqueness, one may investigate what can be said of the elements of the dynamical system. We spend the remainder of this article pursuing these questions. As a preliminary step, we compute the spectrum of the bicrossed product.
Proposition 4.5.
Let $\mathbb{G}$ be a locally compact quantum group and $A$ a $C^{*}$-algebra admitting a faithful state. Assume that there is a coaction $(\mathbb{G},A,\alpha)$ which is Stone-von Neumann unique. Then the spectrum of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ consists of a single point.
Proof.
We begin with a Stone-von Neumann Unique coaction $(\mathbb{G},A,\alpha)$, and must show that any two irreducible representations $\theta,\tilde{\theta}$ of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ on Hilbert spaces $H_{\theta},H_{\tilde{\theta}}$ are equivalent. First, consider $\theta\otimes 1$ as a representation of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ on Hilbert-$A$ module $H_{\theta}\otimes A$, similarly $\tilde{\theta}\otimes 1$. Using the canonical covariant representation of $(\widehat{\mathbb{G}}^{\textnormal{op}},\mathbb{G}\ltimes A,\widehat{\alpha}^{\textnormal{op}})$
$j_{\widehat{\mathbb{G}}^{\textnormal{op}}}\colon C_{0}^{u}(\mathbb{G}^{c})\rightarrow M(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)),\quad j_{\mathbb{G}\ltimes A}\colon\mathbb{G}\ltimes A\rightarrow M(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))$
and then composing with $\theta\otimes 1$ (extended uniquely to a strictly continuous map on the multiplier algebra), we obtain a covariant representation of $(\widehat{\mathbb{G}}^{\textnormal{op}},\mathbb{G}\ltimes A,\widehat{\alpha}^{\textnormal{op}})$ on $H_{\theta}\otimes A$. By Proposition 3.9, we recover the Heisenberg representation $(H_{\theta}\otimes A,\rho^{c}_{\theta},\widehat{\rho}_{\theta},\pi_{\theta})$ of $(\mathbb{G},A,\alpha)$. More concretely,
$\rho^{c}_{\theta}=(\theta\circ j_{\widehat{\mathbb{G}}^{\textnormal{op}}})\otimes 1,\quad\widehat{\rho}_{\theta}=(\theta\circ j_{\mathbb{G}\ltimes A}\circ j_{\mathbb{G}})\otimes 1,\quad\pi_{\theta}=(\theta\circ j_{\mathbb{G}\ltimes A}\circ j_{A})\otimes 1,$
where
$j_{\mathbb{G}}\colon C_{0}^{u}(\widehat{\mathbb{G}})\rightarrow M(\mathbb{G}\ltimes A),\quad j_{A}\colon A\rightarrow M(\mathbb{G}\ltimes A)$
are the canonical maps. We can perform a similar construction for $\tilde{\theta}$. By our assumption, there is a set $S_{\theta}$ so that $(H_{\theta}\otimes A,\rho^{c}_{\theta},\widehat{\rho}_{\theta},\pi_{\theta})$ is a $|S_{\theta}|$-multiple of the Schrödinger representation. Similarly, ${(H_{\tilde{\theta}}\otimes A,\rho^{c}_{\tilde{\theta}},\widehat{\rho}_{\tilde{\theta}},\pi_{\tilde{\theta}})}$ is a Heisenberg representation which is equivalent to a $|S_{\tilde{\theta}}|$-multiple of the Schrödinger representation. Let $S=S_{\theta}\times S_{\tilde{\theta}}\times\mathbb{N}$. By cardinal arithmetic, we have $|S\times S_{\theta}|=|S|=|S\times S_{\tilde{\theta}}|$, hence $|S|$-multiples of $(H_{\theta}\otimes A,\rho^{c}_{\theta},\widehat{\rho}_{\theta},\pi_{\theta})$ and $(H_{\tilde{\theta}}\otimes A,\rho^{c}_{\tilde{\theta}},\widehat{\rho}_{\tilde{\theta}},\pi_{\tilde{\theta}})$ are equivalent. Let
$U\colon\bigoplus_{S}H_{\theta}\otimes A\rightarrow\bigoplus_{S}H_{\tilde{\theta}}\otimes A$
be the corresponding isomorphism which implements this equivalence.
By our assumption, there is a faithful state $\varpi$ on $A$, so let $K_{\varpi}$ be the corresponding GNS Hilbert space and $\Lambda_{\varpi}$ the corresponding GNS map. Let us consider
(4) $\text{id}\otimes\Lambda_{\varpi}\colon H_{\theta}\otimes A\rightarrow H_{\theta}\otimes K_{\varpi}$
which on a linearly dense set acts via $\xi\otimes a\mapsto\xi\otimes\Lambda_{\varpi}(a)$. Take an arbitrary linear combination from this set $\sum_{i}\xi_{i}\otimes a_{i}\in H_{\theta}\odot A$ and calculate
$\begin{split}&\quad\;\bigl{\|}(\text{id}\otimes\Lambda_{\varpi})\bigl{(}\sum_{i}\xi_{i}\otimes a_{i}\bigr{)}\bigr{\|}^{2}=\sum_{i,j}\langle\xi_{i}|\xi_{j}\rangle\langle\Lambda_{\varpi}(a_{i})|\Lambda_{\varpi}(a_{j})\rangle=\sum_{i,j}\langle\xi_{i}|\xi_{j}\rangle\varpi(a_{i}^{*}a_{j})\\ &=\varpi\bigl{(}\bigl{\langle}\sum_{i}\xi_{i}\otimes a_{i},\sum_{i}\xi_{i}\otimes a_{i}\bigr{\rangle}\bigr{)}\leq\bigl{\|}\bigl{\langle}\sum_{i}\xi_{i}\otimes a_{i},\sum_{i}\xi_{i}\otimes a_{i}\bigr{\rangle}\bigr{\|}=\bigl{\|}\sum_{i}\xi_{i}\otimes a_{i}\bigr{\|}^{2}.\end{split}$
It follows that (4) is a well defined contraction. Next we claim that
(5) $\langle(\text{id}\otimes\Lambda_{\varpi})(x)|(\text{id}\otimes\Lambda_{\varpi})(x^{\prime})\rangle=\varpi(\langle x,x^{\prime}\rangle).$
for $x,x^{\prime}\in H_{\theta}\otimes A$. Both sides are sesquilinear and continuous in $x,x^{\prime}$, hence it is enough to check (5) on a linearly dense set: take $x=\xi\otimes a,x^{\prime}=\xi^{\prime}\otimes a^{\prime}$. Then
$\langle(\text{id}\otimes\Lambda_{\varpi})(x)|(\text{id}\otimes\Lambda_{\varpi})(x^{\prime})\rangle=\langle\xi\otimes\Lambda_{\varpi}(a)|\xi^{\prime}\otimes\Lambda_{\varpi}(a^{\prime})\rangle=\langle\xi|\xi^{\prime}\rangle\varpi(a^{*}a^{\prime})=\varpi(\langle x,x^{\prime}\rangle).$
Note that from (5) follows that $\text{id}\otimes\Lambda_{\varpi}$ is injective. A similar reasoning can be performed for $\tilde{\theta}$ – we will abuse the notation and denote the corresponding map also by $\text{id}\otimes\Lambda_{\varpi}$. Next let us define a linear map between Hilbert spaces
$\tilde{U}\colon\bigoplus_{S}H_{\theta}\otimes K_{\varpi}\rightarrow\bigoplus_{S}H_{\tilde{\theta}}\otimes K_{\varpi}.$
First, on a dense subspace $\bigoplus_{S}(\text{id}\otimes\Lambda_{\varpi})(H_{\theta}\odot A)$, we set
(6) $\tilde{U}\bigl{(}((\text{id}\otimes\Lambda_{\varpi})\Omega_{i})_{i\in S}\bigr{)}=\bigl{(}(\text{id}\otimes\Lambda_{\varpi})(\zeta_{i})\bigr{)}_{i\in S}\textnormal{ where }U\bigl{(}(\Omega_{i})_{i\in S}\bigr{)}=(\zeta_{i})_{i\in S}.$
Since $\varpi$ is faithful, $\tilde{U}$ is well defined on a dense subset. Let us argue that it is isometric. Take $((\text{id}\otimes\Lambda_{\varpi})\Omega_{i})_{i\in S}$ and $\zeta_{i}$ as above. Then using (5) we have
$\begin{split}&\quad\;\bigl{\|}\tilde{U}\bigl{(}((\text{id}\otimes\Lambda_{\varpi})\Omega_{i})_{i\in S}\bigr{)}\bigr{\|}^{2}=\bigl{\|}\bigl{(}(\text{id}\otimes\Lambda_{\varpi})(\zeta_{i})\bigr{)}_{i\in S}\bigr{\|}^{2}=\sum_{i\in S}\|(\text{id}\otimes\Lambda_{\varpi})(\zeta_{i})\|^{2}\\ &=\sum_{i\in S}\varpi(\langle\zeta_{i},\zeta_{i}\rangle)=\varpi\bigl{(}\sum_{i\in S}\langle\zeta_{i},\zeta_{i}\rangle\bigr{)}=\varpi\bigl{(}\bigl{\langle}(\zeta_{i})_{i\in S},(\zeta_{i})_{i\in S}\bigr{\rangle}\bigr{)}\\ &=\varpi\bigl{(}\bigl{\langle}U\bigl{(}(\Omega_{i})_{i\in S}\bigr{)},U\bigl{(}(\Omega_{i})_{i\in S}\bigr{)}\bigr{\rangle}\bigr{)}=\varpi\bigl{(}\bigl{\langle}(\Omega_{i})_{i\in S},(\Omega_{i})_{i\in S}\bigr{\rangle}\bigr{)}=\sum_{i\in S}\varpi(\langle\Omega_{i},\Omega_{i}\rangle)\\ &=\sum_{i\in S}\langle(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}|(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}\rangle=\bigl{\|}\bigl{(}(\text{id}\otimes\Lambda_{\varpi})\Omega_{i}\bigr{)}_{i\in S}\bigr{\|}^{2}\end{split}$
This shows that $\tilde{U}$ is an isometry. Clearly it has dense image; consequently, it is unitary.
Next, we check that $\tilde{U}$ is an intertwiner between two representations of $M(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))$, the first being $\bigoplus_{S}\theta\otimes 1$ on $\bigoplus_{S}H_{\theta}\otimes K_{\varpi}$ and the other $\bigoplus_{S}\tilde{\theta}\otimes 1$ on $\bigoplus_{S}H_{\tilde{\theta}}\otimes K_{\varpi}$. From (6) it follows that
$\tilde{U}\bigl{(}\bigoplus_{S}(\text{id}\otimes\Lambda_{\varpi})\bigr{)}=\bigl{(}\bigoplus_{S}(\text{id}\otimes\Lambda_{\varpi})\bigr{)}U\quad\textnormal{ on }\bigoplus_{S}H_{\theta}\otimes A.$
We need to check that
$\tilde{U}\bigl{(}\bigoplus_{S}\theta(x)\otimes 1\bigr{)}=\bigl{(}\bigoplus_{S}\tilde{\theta}(x)\otimes 1\bigr{)}\tilde{U}$
for $x\in M(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))$. However, since $M(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))$ is generated by the images of $C_{0}^{u}(\mathbb{G}^{c}),C_{0}^{u}(\widehat{\mathbb{G}}),A$ it is enough to consider $x$ of the form $j_{\widehat{\mathbb{G}}^{\textnormal{op}}}(y)$, $j_{\mathbb{G}\ltimes A}\circ j_{\mathbb{G}}(z)$ and $j_{\mathbb{G}\ltimes A}\circ j_{A}(a)$. Take $a\in A$.
$\begin{split}&\quad\;\tilde{U}\bigl{(}\bigoplus_{S}\theta(j_{\mathbb{G}\ltimes A}\circ j_{A}(a))\otimes 1\bigr{)}(\xi_{i}\otimes\Lambda_{\varpi}(a_{i}))_{i\in S}=\tilde{U}(\theta(j_{\mathbb{G}\ltimes A}\circ j_{A}(a))\xi_{i}\otimes\Lambda_{\varpi}(a_{i}))_{i\in S}\\ &=\tilde{U}\bigl{(}\bigoplus_{S}(\text{id}\otimes\Lambda_{\varpi})\bigr{)}(\theta(j_{\mathbb{G}\ltimes A}\circ j_{A}(a))\xi_{i}\otimes a_{i})_{i\in S}=\bigl{(}\bigoplus_{S}(\text{id}\otimes\Lambda_{\varpi})\bigr{)}U(\theta(j_{\mathbb{G}\ltimes A}\circ j_{A}(a))\xi_{i}\otimes a_{i})_{i\in S}\\ &=\bigl{(}\bigoplus_{S}(\text{id}\otimes\Lambda_{\varpi})(\tilde{\theta}(j_{\mathbb{G}\ltimes A}\circ j_{A}(a))\otimes 1)\bigr{)}U(\xi_{i}\otimes a_{i})_{i\in S}\\ &=\bigl{(}\bigoplus_{S}\tilde{\theta}(j_{\mathbb{G}\ltimes A}\circ j_{A}(a))\otimes 1\bigr{)}\bigl{(}\bigoplus_{S}(\text{id}\otimes\Lambda_{\varpi})\bigr{)}U(\xi_{i}\otimes a_{i})_{i\in S}\\ &=\bigl{(}\bigoplus_{S}\tilde{\theta}(j_{\mathbb{G}\ltimes A}\circ j_{A}(a))\otimes 1\bigr{)}\tilde{U}(\xi_{i}\otimes\Lambda_{\varpi}(a_{i}))_{i\in S}.\end{split}$
The remaining conditions can be checked in a completely analogous way. This shows that $\bigoplus_{S}\theta\otimes 1$ and $\bigoplus_{S}\tilde{\theta}\otimes 1$ are equivalent. Consequently, there are multiples of $\theta$ and $\tilde{\theta}$ which are equivalent. But since $\theta,\tilde{\theta}$ are assumed to be irreducible, it follows that they are equivalent themselves ([DixmierC, Proposition 5.3.3]). In particular, they are equivalent as representations of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ and consequently this $C^{*}$-algebra has spectrum consisting of one point.
Let us note here that the above proposition applies to any separable $C^{*}$-algebra, since such a $C^{*}$-algebra admits a faithful state.
With this in hand, we will now see that a complete characterization of coefficient algebras $A$ boils down to Naimark’s problem [naimark]. To this end, let us make a general observation that for the trivial coaction $(\mathbb{G},A,\operatorname{triv})$, we have
(7) $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\simeq(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))\otimes_{\textnormal{max}}A.$
Corollary 4.6.
Let $\mathbb{G}$ be a locally compact quantum group and $A$ a separable $C^{*}$-algebra, and assume that $(\mathbb{G},A,\operatorname{triv})$ enjoys Stone-von Neumann uniqueness. Then $A$ is an elementary $C^{*}$-algebra
Proof.
Note that since $A$ is separable, it has a faithful state. Let $\vartheta_{2},\tilde{\vartheta_{2}}$ be irreducible representations of $A$ and $\vartheta_{1}$ an irreducible representation of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})$. Then we can form a representation $\vartheta_{1}\otimes\vartheta_{2}$ of $(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))\otimes_{\textnormal{max}}A$ on $H_{\vartheta_{1}}\otimes H_{\vartheta_{2}}$; same for $\vartheta_{1}\otimes\tilde{\vartheta}_{2}$. They are irreducible, hence by isomorphism (7) and Proposition 4.5, there is a unitary $v\colon H_{\vartheta_{1}}\otimes H_{\vartheta_{2}}\rightarrow H_{\vartheta_{1}}\otimes H_{\tilde{\vartheta}_{2}}$ such that $v(\vartheta_{1}\otimes\vartheta_{2})(\cdot)=(\vartheta_{1}\otimes\tilde{\vartheta}_{2})(\cdot)v$. Extend these representations to multiplier algebras, and consider representations $A\ni a\mapsto(\vartheta_{1}\otimes\vartheta_{2})(1\otimes a)=1\otimes\vartheta_{2}(a)\in\mathbb{B}(H_{\vartheta_{1}}\otimes H_{\vartheta_{2}})$, and same for $\tilde{\vartheta}_{2}$. These are multiples of $\vartheta_{2},\tilde{\vartheta}_{2}$ which are equivalent via the unitary $v$. Since $\vartheta_{2},\tilde{\vartheta}_{2}$ are irreducible, we can use again [DixmierC, Proposition 5.3.3] to deduce that $\vartheta_{2},\tilde{\vartheta}_{2}$ are equivalent. We assume that $A$ is separable, hence [DixmierC, Exercise 4.7.2] tells us that $A$ is elementary.
We also observe that separability is an important obstruction to a general characterization of algebras with one-point spectrum. Akemann and Weaver demonstrate in [awCounterexample] that the existence of nonelementary $C^{*}$-algebras with one-point spectrum and $\aleph_{1}$ generators is independent of ZFC.
Corollary 4.7.
Let $(\mathbb{G},A,\alpha)$ be a left coaction, where $\mathbb{G}$ is a regular locally compact quantum group and $A$ a $C^{*}$-algebra which admits a faithful state. If $(\mathbb{G},A,\alpha)$ is subject to Stone-von Neumann uniqueness, then $\alpha$ is a maximal coaction.
Proof.
We have a canonical surjective ***-homomorphism $\rho\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$, and we want to show that this map is injective. Assume otherwise, i.e. assume that $\text{ker}(\rho)$ is a nontrivial ideal in $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$. Then we can form quotient $C^{*}$-algebra $(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))/\text{ker}(\rho)$ and choose its irreducible representation $\theta\colon(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))/\text{ker}(\rho)\rightarrow\mathbb{B}(H_{\theta})$. Composing with the quotient map, we obtain an irreducible representation $\tilde{\theta}\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{B}(H_{\theta})$ such that $\text{ker}(\rho)\subseteq\text{ker}(\tilde{\theta})$. On the other hand, since $\text{ker}(\rho)$ is nontrivial, we can choose $x\in\text{ker}(\rho)\setminus\{0\}$ and irreducible representation $\tilde{\theta}^{\prime}\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{B}(H_{\tilde{\theta}^{\prime}})$ such that $\tilde{\theta}^{\prime}(x)\neq 0$. This shows that $\tilde{\theta},\tilde{\theta}^{\prime}$ have different kernels, in particular are not equivalent – this gives us a contradiction with Proposition 4.5. Thus $\rho$ is an isomorphism, and the coaction $(\mathbb{G},A,\alpha)$ is maximal.
As a final application, we identify another sufficient condition for strong regularity of quantum group $\widehat{\mathbb{G}}$.
Corollary 4.8.
Let $\mathbb{G}$ be a regular locally compact quantum group and assume that $(\mathbb{G},\mathbb{C},\operatorname{triv})$ is Stone-von Neumann unique. Then $\widehat{\mathbb{G}}$ is strongly regular.
Proof.
The previous corollary gives us an isomorphism $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))$. Let $j_{{\widehat{\mathbb{G}}}^{\textnormal{op}}}\colon C_{0}^{u}(\mathbb{G}^{c})\rightarrow M({\widehat{\mathbb{G}}}^{\textnormal{op}}\ltimes C_{0}(\widehat{\mathbb{G}}))$ be the canonical inclusion. Observe that $(j_{{\widehat{\mathbb{G}}}^{\textnormal{op}}},j_{C_{0}(\widehat{\mathbb{G}})}\hat{\lambda})$ is a covariant representation of the system $(\widehat{\mathbb{G}}^{\textnormal{op}},C_{0}^{u}(\widehat{\mathbb{G}}),(\hat{\lambda}\otimes\text{id})\circ\Delta^{u}_{\widehat{\mathbb{G}}^{\textnormal{op}}})$ on $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}(\widehat{\mathbb{G}})$ whose integrated form determines a (nonzero) surjection of $\mathbb{K}(L^{2}(\mathbb{G}))$. Since $\mathbb{K}(L^{2}(\mathbb{G}))$ is simple, it has to be an isomorphism. We conclude that $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}(\widehat{\mathbb{G}})$ comprises compact operators so that by definition, $\widehat{\mathbb{G}}^{\textnormal{op}}$ is strongly regular. This property is equivalent to strong regularity of $\widehat{\mathbb{G}}$, which ends the proof. ∎